Thursday, March 31, 2011

Popping segments in & swapping RSS bits

Janus & Luis are still working on the guidance system - both in the hardware & software departments.  This is a critical aspect of the telescope's functionality & something we'd really like to get going before RSS goes up next week.

In the meantime, Jonathan, Hitesh & Denville went about replacing one of the mirror mounts that'd suffered a broken flexure.

One gets a sense of the relative ages of the mirror coatings in this low-angle view of the segments.  It really isn't nearly as bad as this looks, a camera flash at this angle is particularly unflattering to even the cleanest of surfaces...

These mirror installations call for incredibly careful cherry-picker driving & Jonathan's got it taped!

Hands reach up from below the truss & guide the mirror into place.

Then 3 threaded rods attached to the mount from below prevent any clocking of the segment while it's being lowered into place..

Here are the cleanest segments along the bottom edge of the array.

Sinking into position...

The primary's way too much fun to photograph!

Very cool (almost) symmetries...

Once in position, a segement can be aligned to within about 50 arcsec in tip/tilt, just by matching it to its neighbours as viewed from the catwalk.  That's enough to get it within the course alignment capture range for the CCAS instrument.  The piston (Z) needs to be set with a spherometer which measures the height of the mirror with respect to those around it, then a full CCAS alignment takes care of the rest.  Individual segments are aligned in tip & tilt to about 0.1 arcsec & the RMS for the whole array comes out to no more than 0.3 arcsec RMS.

Over on RSS it was time to test the changing of the slitmask magazine, the filter magazine & the Fabry-Perot etalons.  None of them trivial operations, of course, but these guys do enjoy a challenge...

Prof Peter keeps an eye while Keith tries an etalon change.  The etalon handling jig (the contraption hanging from the cable ) isn't Quite ready so for now it's still easier to do this by hand, but the planned mods to the jig will make this a safer & simpler operation - especially when doing this up on the tracker!

Nightlog 2011-03-29

SA: David
SO: Siphelo
Others: Darragh, Luis, Deneys, Janus, Eben


Clouds and then rain - no observing.

Further SW work by Luis & Janus to get guiding ready.


Tuesday, March 29, 2011

Second Light for SALTICAM!

Yesterday was a classic Blue Monday with various things going wrong - from an O-ring popping out of one of the Cryotiger hoses & sending our CCD cooling efforts back to square one, to a potential train-smash with the SALTICAM auto-guider...

The guys did, however, manage to baffle the whole payload, connect SALTICAM up to everything, establish comms between it & the control room, fix some RSS hardware issues & sort out the glycol for the igloo that will house the 2 Cryotigers.

Then, by about 9pm, we were authorised to proceed with SALTICAM tests, even though the CCDs weren't cooled & the dome hadn't been conditioned.  So there was nothing pretty about the images, or the stack, but some good progress was made in verifying instrument & software functionality none the less.

Today's been a *much* better day.  The best news of the morning was finding that the fibres from SALTICAM's auto-guider are ok afterall - a Huge relief!  Later Darragh & Luis talked to everyone about the workings of SALTICAM & its software, including background about all the tests to be done & their relevance in the grand commissioning scheme of things.

The good news for the evening was that the CCDs had cooled down as required & that the weather was Perfect, so the stage was set for SALTICAM's proper Second Light...

The substantial crowd gathered in the control room was treated to beautiful images of the globular cluster NGC 2808 - uniform stars over the entire field of view :)

Happy Day indeed & great to have so many people around to share it with!

BVRI images were taken & Petri combined them to produce this wonderful BVR colour version of the cluster.

The crowd slowly dispersed, but the show went on with Petri & Zolisa running the telescope while Darragh & Luis worked through their list of tests & software fixes.

The novelty of consistently getting such stunning images seems unlikely to wear off any time soon! :) :) :)

Nightlog 2011-03-28

SA: Petri
SO: Zolisa
Others: Darragh, Luis, Deneys, Janus, Lisa, Ockert, ...


- The rotating structure and Salticam were lifted and aligned
on Sunday. The ccd was cooled monday, but there was a hitch
and it still needs to continue on tuesday.

- Nevertheless, went on-sky with a warm ccd and a very
un-conditioned dome to check basic functionalities of
Salticam, e.g. slot-mode and FT work, filter cycles etc.

Salticam second light!!


Sunday, March 27, 2011

The payload goes back up!

Further trial-fitting of the ADC showed that the installation, to be done in a couple of weeks time with everything else in place up there, would be way too difficult with such limited access & minimal clearance.  Yet more nerve-wracking NRS surgery was needed so we had Jonathan on the vacuum cleaner & Eben driving the router - just a few centimetres above the SAC!

An elaborate scheme to keep all the carbon fibre shavings from landing up on M5 & M3 involved a sheet of bubble wrap, a piece of corex, adhesive covering material from Eben's model airplanes (sticky side up), plenty of duct tape, a vacuum cleaner & some good karma.  We were all Very glad when this process was over!

Next on our Sunday afternoon line-up was the lifting of the payload...  With all the hatches bolted on, the rotating structure was wheeled out into the loading bay.

The whole lot was rigged up & tilted to the required 37 degrees before being lifted with the dome crane.

Up through the hatch leading to the telescope floor.

& then onward & upward to the tracker...

Charl, Eben, Jono & Chris were up there to meet it.

& guided it down into position inside the NRS.

Once the payload was bolted into position & the team felt sure that everything was clear, they rotated rho to make sure nothing crashed or pinched.

No problems - well done guys!  Nice of Eben & Charl to coordinate their outfits for the big occasion, don't you think?

Prior to the lift, the SAC reference mirror was bolted onto the back of M4 & then after dinner we went up the CCAS tower to check the alignment of the payload with respect to the SAC. 

With the tracker at its static offsets, the alignment telescope in the tower was trained on the SAC reference mirror's cross-hair & then squared up in tip & tilt to get the alignment telescope on the optical axis.

A target for the alignment telescope was mounted in a fixture that centrally locates in the floor of the payload & then the bar with the cross-hair (used in the SALTICAM alignment) was bolted in place on the top of the rotating structure - the 2 then define the axis of the payload.

Although not a hugely sensitive test, since the separation between these 2 points is small compared to their distance from the alignment telescope, the target & cross-hair were well aligned.  This shows that the payload tip/tilt is ok & since this point was close enough to the optical axis (given the known problem of the rho stage wobble), no adjustments were made to the rotating structure.

It was, in fact, time to put an end to this extremely busy weekend!  Down the hill, to the porch behind the hostel where, under an absolutely perfect Karoo sky & midst a mini-plague of locusts, we toasted SALTICAM's return to the tracker :)

Saturday, March 26, 2011

Getting to know RSS

Today Ant & I worked through the acceptance test procedure (ATP) for his RSS software.  This involved testing all the mechanisms & making sure that everything did as it was told & that the software reported each event accordingly.  We also deliberately shut down the software mid-way through various moves & procedures, to see what would happen & whether the system would know where everything was when it came back up.  

Other than the temperamental waveplates, everything went extremely smoothly & the only real trouble came when we attempted to recover from the trickiest of all scenarios: the aftermath of clicking the "kill power" button while everything was moving.  This is a Very big ask for any system, particularly one as complex as RSS & so it's actually pretty impressive that only the slitmask mechanism got itself really confused & couldn't be consoled by being homed.  Luckily Vic was around to coax it back mechanically so it's all good!

Activating all the mechanisms provided a great photo-op, which permits an introduction to the various bits & pieces that make up this Swiss Army Knife spectrograph.  The quarter & half waveplates were the only mechanisms not photographed, since they're hidden by their baffled-up housing.  First up - the shutter, closed & open.

Then the grating holder - empty & with a grating in place, lowered in from the grating magazine above.  The magazine holds 6 transmission gratings, 5 of them of the volume phase holographic (VPH) kind & another of the more traditional surface relief sort.  The latter has a line frequency of 300 lines/mm while the VPH gratings have 900, 1300, 1800, 2300 & 3000 lines/mm.

A characteristic of VPH gratings  is that their efficiency varies with input angle & so a single grating can cover a large wavelength range with good efficiency if one can change the relative angle between the collimated beam and the grating normal.  The highest resolution VPH grating (3000 lines/mm) is shown below, at angles of 20, 30, 40 & 50 degrees to the beam.

The image below is of the low resolution, surface relief grating at angles of zero & 50 degrees to the beam.

The RSS camera articulates in order to preserve the Littrow configuration (in which the camera/collimator angle is twice that of the light incident on the grating) since this maximises the efficiency of the grating.  Below you see the camera at articulation angles of zero & 70 degrees to the beam.

The beamsplitter needs to be inserted in order to do any kind of polarimetry.  This fascinating optic was described in detail in an earlier post, but here we show it in its 2 positions, out of the beam & in the beam.

RSS employs a number of filters: 5 for order-blocking (to suppress light arising from higher order diffraction) & more than 30 interference filters that are used in Fabry-Perot mode.  A filter magazine travels along the outside of the camera & once the chosen filter station's reached, the filter gets pneumatically inserted into the beam.  Here we see the magazine at stations 1 & 10 of 20.  Filters will need to be changed periodically based on the observing schedule.

Then there are the Fabry-Perot etalons, which are also pneumatically inserted into & removed from the beam.  Only 2 of the 3 will be in the instrument at at given time - always the low resolution one in conjunction with either the medium or high resolution etalon.

The slimasks (to be used in long-slit & multi-object spectroscopy modes) are also stored within a magazine, but space constraints don't allow for a traveling magazine.  Instead there is an elevator that travels up & down to fetch the desired mask & bring it to the slit plane.

The mask is then pneumatically inserted or removed.  The RSS auto-guider attaches below the slit-plane & is not shown in the pair of images below.

All the while that we were playing around with RSS, Charl diligently worked away at tidying & baffling up the payload which we're hoping to lift tomorrow...